# Phenomenology of a singlet-doublet-triplet scotogenic framework

### v1

@@ 4.2 Dark matter
4.3 Lepton flavour violation
@@
Table 2: Constraints from the Higgs-boson mass ($m_{H}$), the DM relic density ($\Omega_{\rm CDM}h^{2}$), and numerous LFV observables (remaining entries) ParticleDataGroup:2022pth as implemented in our MCMC analysis.
@@
In our MCMC study, we include all available observables related to lepton flavour violating transitions, as summarised in Table 2. We make use of SPheno 4.0.5 to numerically compute the associated predicted values for a given parameter point.
@@ 5.2 Dark matter mass and phenomenology
Note that we do not take into account constraint stemming from indirect detection, as their realistic numerical implementation is quite heavy in a framework involving several additional fields as concluded in Ref. Armand:2022sjf .

### v3

@@ 4.2 Dark matter
4.3 Lepton flavour and $CP$ violation
@@
$\text{EDM}_{e}$ $<1.1\times 10^{-27}$
$\text{EDM}_{\mu}$ $<1.8\times 10^{-19}$
$\text{EDM}_{\tau}$ $<1.15\times 10^{-15}$
Table 2: Constraints from the Higgs-boson mass ($m_{H}$), the DM relic density ($\Omega_{\rm CDM}h^{2}$), and numerous LFV and $CP$-violating observables (remaining entries) ParticleDataGroup:2022pth as implemented in our MCMC analysis.
@@
In our MCMC study, we include all available observables related to lepton flavour violating transitions, as summarised in Table 2. In addition, we include the constraints on the electric dipole moments of the electron, muon, and tau lepton, as indicated in Table 2 ParticleDataGroup:2022pth . We make use of SPheno 4.0.5 to numerically compute the associated predicted values for a given parameter point. Let us already note that the electric dipole moments (EDMs) do not have a significant impact within our study as all additional $CP$-violating effects with respect to the Standard Model turn out to be negligible.
@@ 5.2 Dark matter mass and phenomenology
It should be noted that the Sommerfeld effect has not been included in the present analysis. For scenarios where the masses of charged particles are close to the DM mass, the Sommerfeld effect could become relevant. However, as shown in Ref. Beneke:2020vff, the resulting modification to the relic density is not expected to alter the final conclusions. Incorporating Sommerfeld factors into the computation of the relic density may lead to a shift in the peak observed in Fig. 6, but this would not impact the overall conclusions of this work.
@@
Note that our analysis does not include constraints stemming from indirect detection. We have, however, verified that the corresponding cross-sections remain below the corresponding limits given in Ref. Fermi-LAT:2016uux. In conclusion, the indirect detection of dark matter does not affect the Likelihood and the results in this specific study. Note also that the realistic numerical implementation of indirect detection constraints in a framework involving several additional fields is quite heavy as concluded in Ref. Armand:2022sjf .
@@ 5.4 Comment on $(g-2)_{\mu}$ and EDM
A related observable to $(g-2)$ is the electric dipole moment (EDM). While the Standard Model contributions for the lepton EDMs are CKM-suppressed, new physics contributions are constrained by the upper limits given in Table 2. In the present setup, as already mentioned above, the smallness of the involved couplings leads to a suppression of additional contributions to the lepton EDMs. The impact of these constraints related to $CP$-violation remain therefore negligible.